# Boundary Control and Calderón type Inverse Problems in Non-local heat equation

### v2

@@ 3. Calderón type Inverse Problems:
$-\partial_{t}(w_{1}-w_{2})+(-\Delta)^{a}(w_{1}-w_{2})+&(q_{1}w_{1}-q_{2}w_{2})=0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ (w_{1}-w_{2})=&0\qquad\qquad\ \ \quad\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ (w_{1}-w_{2})(T,\cdot)=&0\qquad\qquad\qquad\qquad\quad\ \ \qquad\mbox{in}\ \Omega,\end{aligned}\right.$
Thanks to (8), (10), (11), we can apply Pohozaev identity (12). It yields
$\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2})(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}$ $(w_{1}-w_{2})(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{w_{1}-w_{2}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
Equation corresponding to $w_{1}-w_{2}$ yields:
(31) $\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))$ $(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x$
-- $\frac{2a-n}{2}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{(w_{1}-w_{2})}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
We intend to calculate the term ${\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x}$. Gauss divergence theorem yields:
$\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))$ $(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=-\int_{\Omega}\nabla\cdot\left(x(-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x$
$=$ $-\int_{\Omega}n(-\Delta)^{a}(w_{1}-w_{2})(w_{1}-w_{2})\,\rm{d}x-\int_{\Omega}x\cdot\nabla\left((-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x.$
Again equation corresponding to $w_{1}-w_{2}$ yields
(32) $\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=$ $-n\int_{\Omega}\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x+n\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x$
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x.$
We compute the last term of the right hand side above:
$-\int_{\mathbb{R}^{N}}x\cdot\nabla$ $\left((-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x=-\int_{\mathbb{R}^{N}}\nabla_{\xi}\cdot\left(\xi|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})\right)(\hat{w_{1}}-\hat{w_{2}})\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}\left(n|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})+2a|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\hat{w_{1}}-\hat{w_{2}}\right)\right)(\hat{w_{1}}-\hat{w_{2}})\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}n|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}+2a|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\frac{(\hat{w_{1}}-\hat{w_{2}})^{2}}{2}\right)\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}(n+2a)|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}\,\rm{d}\xi+\int_{\mathbb{R}^{N}}\nabla_{\xi}.\left(\xi|\xi|^{2a}\right)\frac{(\hat{w_{1}}-\hat{w_{2}})^{2}}{2}\,\rm{d}\xi$
$=$ $-\frac{n+2a}{2}\int_{\mathbb{R}^{N}}|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}\,\rm{d}\xi=-\frac{n+2a}{2}\int_{\Omega}(-\Delta)^{a}(w_{1}-w_{2})(w_{1}-w_{2})\,\rm{d}x$
Replacing $(-\Delta)^{a}(w_{1}-w_{2})$ in terms of time derivative yields
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}(w_{1}-w_{2})\right)$ $(w_{1}-w_{2})\,\rm{d}x$
@@
$=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}(w_{1}-w_{2})(w_{1}-w_{2})-(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\right)\,\rm{d}x$
$=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)-(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\right)\,\rm{d}x.$
Substituting this in (32), we obtain
$\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=-\frac{3n+2a}{2}\int_{\Omega}$ $\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x$
$+\frac{3n+2a}{2}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x.$
Substituting this in (31) yields
$(n+2a)\int_{\Omega}\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x-(n+2a)\int_{\Omega}$ $(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{(w_{1}-w_{2})}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma=0.$
Integrating with respect to time yields:
$\int_{\Omega}(w_{1}-w_{2})^{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}(w_{1}-w_{2})^{2}(t_{2},\cdot)$ $\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})\,\rm{d}x\rm{d}t$
$=$ $-\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{(w_{1}-w_{2})}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}t.$
The equality (30) yields:
(33) $\int_{\Omega}(w_{1}-w_{2})^{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}(w_{1}-w_{2})^{2}(t_{2},\cdot)$ $\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})\,\rm{d}x\,\rm{d}t=0$
We recall equation corresponding to $w_{1}-w_{2}$:
$-\partial_{t}(w_{1}-w_{2})+(-\Delta)^{a}(w_{1}-w_{2})+(q_{1}w_{1}-q_{2}w_{2})=0$
with $w_{1}-w_{2}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$. We multiply the equation corresponding to $(w_{1}-w_{2})$ by $(w_{1}-w_{2})$ and integrate with respect to space and time.
$\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{2},\cdot)\,\rm{d}x$ $+\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}(w_{1}-w_{2})(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t=0.$
@@
$\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{2},\cdot)\,\rm{d}x+$ $\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}(w_{1}-w_{2})\right|^{2}\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t=0,$
@@
$\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{2},\cdot)\,\rm{d}x+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t$
$=$ $-\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}(w_{1}-w_{2})\right|^{2}\,\rm{d}x\,\rm{d}t=0.$
By Hardy-Littlewood-Sobolev estimate:
$\|(w_{1}-w_{2})\|_{\mathrm{L}^{2}(\Omega)}\leq C_{Vol}\|(w_{1}-w_{2})\|_{\mathrm{L}^{\frac{2n}{n-a}}(\Omega)}\leq C_{n,a,\Omega}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}(w_{1}-w_{2})\right|^{2}\,\rm{d}x,$
where $C_{Vol}$ and $C_{n,a,\Omega}$ are two fixed positive constants depends on the domain and the dimension only. Hence
$\|w_{1}-w_{2}\|_{\mathrm{L}^{2}((t_{1},t_{2})\times\Omega)}=0,\quad\forall\,t_{1}\leq t_{2}\in[0,T].$
@@
It further implies $(q_{1}-q_{2})w_{1}\equiv 0$. Let’s fix a time $t\in[0,T]$. We have
$\int_{\Omega}(q_{1}-q_{2})(t,\cdot)w_{1}(t,\cdot)\,\rm{d}x=0.$
Thanks to theorem 2, we have that $w_{1}(t,\cdot)$ is dense in $\mathrm{L}^{2}(\Omega)$. Hence $q_{1}(t,\cdot)\equiv q_{2}(t,\cdot)$. Since $t$ is arbitrary, we obtain
@@ 4. Further result:
(34) {$(-\Delta)^{a}\phi=&\lambda\phi\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ \phi=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\Omega.\end{aligned}\right.$
@@
The following regularity of the eigenfunctions can be found in [ROS14a], [FRRO16]:
@@
While we were uploading our work to arXiv, Umberto Biccari, Mahamadi Warma, and Enrique Zuazua also posted their paper titled "Boundary Observation and Control for Fractional Heat and Wave Equations." In their work, they require the fractional power ${a\in\left(\frac{1}{2},1\right)}$ to achieve null controllability.
Acknowledgment: The author was funded by the Department of Atomic Energy ((((DAE)))), Government of India, in the form of Postdoctoral Research Fellowship.

### v3

@@ 3. Calderón type Inverse Problems:
$\left\{\begin{aligned} \partial_{t}V+(-\Delta)^{a}V=&0\qquad\mbox{in}\ (0,T)\times\Omega\\ V=&0\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\overline{\Omega}\\ \frac{V}{d^{a-1}}=&F\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ V(0,\cdot)=&0\qquad\mbox{in}\ \{0\}\times\Omega,\end{aligned}\right$.
where $F(0,\cdot)=0$ and $F\in\mathrm{L}^{2}((0,T);H^{a+\frac{1}{2}}(\partial\Omega))$. Hence Thanks to theorem 8 [Gru23], we have that
$V\in\mathrm{L}^{2}\left((0,T);H^{(a-1)(2a)}(\overline{\Omega})\right)\cap\overline{H}^{1}\left((0,T);\mathrm{L}^{2}(\Omega)\right).$
Integration by parts (16) yields
$\int_{\Omega}V(s,x)(-\Delta)^{a}v(t,x)\,\rm{d}x$- $\int_{\Omega}v(t,x)(-\Delta)^{a}V(t,x)\,\rm{d}x$
$=-\Gamma(a)\Gamma(a+1)\int_{\partial\Omega}\frac{V}{d^{a-1}}(t,\sigma)\frac{v}{d^{a}}(s,\sigma)\,\rm{d}\sigma=0.$
Integrating with respect to variable ‘$t$’ yields
$\int_{0}^{T}\int_{\Omega}V(s,x)(-\Delta)^{a}v(t,x)-v(t,x)(-\Delta)^{a}V(t,x)\,\rm{d}x\,\rm{d}t=0$
Equation (3) and equation (3) yields
$\int_{0}^{T}\int_{\Omega}\left(-\partial_{t}v(t,x)-(q_{1}w_{1}-q_{2}w_{2})(t,x)\right)V(s,x)\,\rm{d}x\,\rm{d}t=\int_{0}^{T}\int_{\Omega}\left(-\partial_{s}V(s,x)\right)v(t,x)\,\rm{d}x\,\rm{d}t.$
Simplifying, it implies
(27) $-\int_{\Omega}v(T,x)V(s,x)\,\rm{d}x+$ $\int_{0}^{T}\int_{\Omega}v(t,x)\partial_{s}V(s,x)\,\rm{d}x\,\rm{d}t$
$=$ $\int_{0}^{T}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(t,x)V(s,x)\,\rm{d}x\,\rm{d}t\quad\mbox{for all}\ s>0.$
Consider the function $w:(0,+\infty)\times(0,+\infty)\to\mathbb{R}$, denoted by
$w(t,s)=\int_{\Omega}v(t,x)V(s,x).$
We will show ‘$w$’ satisfy a particular transport equation. We compute the term $\partial_{t}w-\partial_{s}w$.
$\partial_{t}w-\partial_{s}w=\int_{\Omega}\partial_{t}v(t,x)$ $V(s,x)\,\rm{d}x-\int_{\Omega}v(t,x)\partial_{s}V(s,x)\,\rm{d}x$
$=-\int_{\Omega}V(s,x)$ $(-\Delta)^{a}v(t,x)\,\rm{d}x+\int_{\Omega}v(t,x)(-\Delta)^{a}V(s,x)\,\rm{d}x$
- $\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(t,x)V(s,x)\,\rm{d}x$
Employing integration by parts formula (16), we obtain
$\partial_{t}w-\partial_{s}w=-\int_{\Omega}$ $(q_{1}w_{1}-q_{2}w_{2})(t,x)V(s,x)\,\rm{d}x$
$+$ $\Gamma(a)\Gamma(a+1)\int_{\partial\Omega}\frac{V}{d^{a-1}}(s,\sigma)\frac{v}{d^{a}}(t,\sigma)\,\rm{d}\sigma.$
As ${\frac{v}{d^{a}}\Big|_{\partial\Omega}=0}$, we obtain
$\partial_{t}w-\partial_{s}w=-\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(t,x)V(s,x)\,\rm{d}x.$
Hence ‘$w$’ satisfy the following transport equation
$\left\{\begin{aligned} \partial_{t}w-\partial_{s}w=&-\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(t,x)V(s,x)\,\rm{d}x\qquad\mbox{in}\ (0,+\infty)\times(0,+\infty)\\ w(0,s)=&0\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \qquad\mbox{on}\ \{0\}\times(0,+\infty)\\ w(t,0)=&0\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \qquad\mbox{on}\ (0,+\infty)\times\{0\}.\end{aligned}\right$.
We perform a change of variable $t-s\to\xi$’ and $s\to s$’. We can rewrite the equation corresponding to ‘$w$’ as
$\partial_{\xi}w=-\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(\xi+s,x)V(s,x)\,\rm{d}x.$
Hence for any $s\geq 0$, ‘$w$’ satisfies the following ordinary differential equation:
@@
$\left\{\begin{aligned} \partial_{\xi}w=&-\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(\xi+s,x)V(s,x)\,\rm{d}x\qquad\mbox{in}\ (-s,+\infty)\\ w(-s)=&0.\end{aligned}\right$.
@@
$w(\xi^{\prime},s)=-\int_{-s}^{\xi^{\prime}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(\xi+s,x)V(s,x)\,\rm{d}x\,\rm{d}t.$
Let $\xi^{\prime}+s=T$, then change of variable yields
$w(T,s)=$ $-\int_{0}^{T}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(t,x)V(s,x)\,\rm{d}x\,\rm{d}t$
$=$ $\int_{\Omega}v(T,x)V(s,x)\,\rm{d}x\qquad\mbox{for all}\ (T,s)\in[0,+\infty)\times[0,+\infty).$
@@
$\int_{0}^{T}\int_{\Omega}v(t,x)\partial_{s}V(s,x)\,\rm{d}x\,\rm{d}t=0.$
Integrating the above relation with respect to ‘$s$’ variable, we obtain
$\int_{0}^{S}\int_{0}^{T}\int_{\Omega}v(t,x)\partial_{s}V(s,x)\,\rm{d}x\,\rm{d}t\,\rm{d}s=0.$
Hence
$\int_{0}^{T}\int_{\Omega}v(t,x)V(S,x)\,\rm{d}x\,\rm{d}t=0\qquad\mbox{for all}\ S\geq 0.$
We have $v\in\mathrm{L}^{2}\left((0,T);H^{a}(\overline{\Omega})\right)\subset\mathrm{L}^{2}((0,T);\mathrm{L}^{2}(\Omega))$. Hence the function ${\int_{0}^{T}v(t,x)\,\rm{d}t\in\mathrm{L}^{2}(\Omega)}$ and
$\int_{\Omega}V(S,x)\left(\int_{0}^{T}v(t,x)\,\rm{d}t\right)\,\rm{d}x=0,\quad\mbox{for all}\ S\geq 0.$
We fix a particular $S>0$. then If we vary $F$, the set $\big\{V(S,x):F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\big\}$ is dense in $\mathrm{L}^{2}(\Omega)$ [see Theorem 3]. Hence
$\int_{0}^{T}v(t,x)\,\rm{d}t=0,\qquad\mbox{for all}\,T\geq 0.$
We fix a particular time $t=\tilde{T}>0$. We have
$\int_{0}^{\tilde{T}+h}v(t,x)\,\rm{d}t-\int_{0}^{\tilde{T}-h}v(t,x)\,\rm{d}t=0,$
where $h$ is sufficiently small. It yields
$\frac{1}{2h}\int_{\tilde{T}-h}^{\tilde{T}+h}v(t,x)\,\rm{d}t=0.$
Employing Lebesgue differentiation theorem, we obtain
[…]
@@
From the equation (5), we have that
$(q_{1}-q_{2})w_{1}\equiv 0.$
We fix $t=\tilde{T}>0$, we have that
$\int_{\Omega}(q_{1}-q_{2})(x)w_{1}(\tilde{T},x)\,\rm{d}x=0.$
Again, if we vary $f\in C_{c}^{\infty}((0,T)\times\Omega)$, the set ${\big\{w_{1}(\tilde{T},x):f\in C_{c}^{\infty}((0,T)\times\partial\Omega)\big\}}$ is dense in $\mathrm{L}^{2}(\Omega)$ [see Theorem 2]. Hence
$(q_{1}-q_{2})(\tilde{T},x)=0,\qquad\mbox{for a.e.}\ x\in\Omega.$
Since $\tilde{T}>0$ is arbitrary, we obtain
@@ 4. Further result:
(28) $\left\{\begin{aligned} (-\Delta)^{a}\phi=&\lambda\phi\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ \phi=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\Omega.\end{aligned}\right$.
@@
The following regularity of the eigenfunctions can be found in [ROS14a], [FRRO16].
@@
Acknowledgment: The author was funded by the Department of Atomic Energy ((DAE)), Government of India, in the form of Postdoctoral Research Fellowship.